### v1

@@ 4.2 Design Principle
For the TPC, each hit provides both 2D positional information and timing data, which together can be transformed into a 3D spatial coordinate. Consequently, a TPC track can be naturally represented as a point cloud. Inspired by the PT, a U-Net-based hierarchical architecture is adopted as the backbone. Furthermore, since the hits are distributed irregularly in 3D space, the point clouds are represented as graphs, where the attention mechanism from the original PT is incorporated into the node aggregation process to enhance feature learning.
For the dNN/dxx reconstruction, each TPC hit is labeled either as positive or negative according to the number of primary electrons it contains, as determined from the MC truth information. To streamline the reconstruction, the two-step procedure described in ref. tian2025cluster is unified into a single model. In this formulation, the reconstruction task can be defined as either a node classification problem or, equivalently, a point cloud segmentation problem.
@@ 4.4 Transformer Layer
The core component of the GraphPT architecture is the transformer layer. In the context of GNNs, the transformer layer performs message passing among nodes: each node receives information from its neighbors and aggregates it to update its own representation. The message-passing process can be formalized as:
@@
where $i$ denotes the index of the aggregation node with feature vector $\mathbf{x_{i}}$ and $\mathcal{N}(i)$ represents its neighboring nodes. The function $\beta$ maps the original feature vectors into a latent space, while the function $\alpha$ computes the attention weights used to aggregate neighbor information. The function $\alpha$ is defined as:
@@
The dot-product operator, which is the fundamental mathematical component underlying self-attention, is further examined. It is defined as:

### v2

@@ 4.2 Design Principle
For the TPC, each hit provides both 2D positional information and timing data, which together can be transformed into a 3D spatial coordinate. Consequently, a TPC track can be naturally represented as a point cloud. Inspired by the PT, we adopt a U-Net-based hierarchical architecture as the backbone. In this work, we further extend the framework by representing the irregularly distributed hits in 3D space as graphs, where the attention mechanism from the original PT is incorporated into the node aggregation process. This integration of GNNs allows more effective feature learning within the point cloud representation.
For the dNN/dxx reconstruction, each TPC hit is labeled either as positive or negative according to the number of primary electrons it contains, as determined from the MC truth information. Building on previous approaches, we streamline the reconstruction by unifying the two-step procedure described in ref. tian2025cluster into a single model. In this formulation, the reconstruction task can be defined as either a node classification problem or, equivalently, a point cloud segmentation problem.
@@ 4.4 Transformer Layer
The core component of the GraphPT architecture is the transformer layer. In the context of GNNs, the transformer layer performs message-passing mechanism: each node aggregates information from its neighbors to update its own representation. Formally, the message-passing can be expressed as:
@@
where $i$ denotes the index of the aggregation node with feature vector $\mathbf{x_{i}}$ and $\mathcal{N}(i)$ represents its neighboring nodes. The function $\beta$ maps the original feature vectors into a latent space, while the function $\alpha$ computes the attention weights used to aggregate neighbor information. The attention weights are defined as:
@@
In contrast, the second operator is proposed in this work as a novel alternative. Specifically, the dot-product operator, which constitutes the fundamental mathematical component underlying self-attention, is adapted to the dNN/dxx reconstruction. It is defined as: